# AMUN: Adversarial Machine UNlearning

### v2

@@ 6.2.2 Adversarially Robust Models
We evaluate the effectiveness of AMUN when the trained model is adversarially robust. One of the most effective methods in designing robust models is adversarial training which targets smoothing the model’s prediction function around the training samples Salman et al. (2019). This has been shown to provably enhance the adversarial robustness of the model Cohen et al. (2019). One of the effective adversarial training methods is by using TRADES loss introduced by (Zhang et al., 2019). We will use adversarially trained ResNet-18 models for unlearning $10\%$ of the samples in CIFAR-10. In addition, we will use another defense mechanism that is less costly and more practical for larger models. There is a separate line of work that try to achieve the same smoothness in model’s prediction boundary by controlling the Lipschitz constant of the models Szegedy (2013). The method proposed by Boroojeny et al. (2024) is much faster than adversarial training and their results show a significant improvement in the robust accuracy. We use their clipping method to evaluate the effectiveness of AMUN for unlearning $10\%$ and $50\%$ of the samples from robust ResNet-18 models trained on CIFAR-10.
Results: Table 5 in Appendix E.2 shows the results for the adversarially trained models. For the models with controlled Lipschitz continuity, the results are shows in Table 3 (no access to $\mathcal{D_{\text{R}}}$) and Table 4 in Appendix E.2 (with access to $\mathcal{D_{\text{R}}}$). As the results show, even when there is no access to $\mathcal{D_{\text{R}}}$, AMUN still results in effective unlearning for adversarially robust models; RMIA does not do better than random guessing for predicting $\mathcal{D_{\text{F}}}$ from $\mathcal{D_{\text{T}}}$. As Fig. 3 (right) shows, in the robust models, more than $97\%$ of the adversarial examples are further away from their corresponding training samples, compared to this distance for the original models. However, this does not interfere with the performance of AMUN because these robust models are smoother and tend to be more regularized. This regularization, which prevents them from overfitting to the training samples is in fact a contributing factor to the improved generalization bounds for these models Bartlett et al. (2017). This in itself contributes to enhanced resilience against MIAs. As seen in Tables 3 and 4, even for the clipped models retrained on $\mathcal{D_{\text{R}}}$, the AUC score of RMIA for predicting $\mathcal{D_{\text{R}}}$ from $\mathcal{D_{\text{F}}}$ (FR AUC) is very low, which shows that these smoother models .
@@ 6.3 Continuous Unlearning
We evaluate the performance of the unlearning methods when they are used to perform multiple consecutive unlearning from a trained model. This is a desirable capability for unlearning methods because in real-world applications there might be multiple unlearning requests and it is preferred to minimize the number of times that a model needs to be retrained from scratch. The setting we envision is as follows: models are updated at each request for unlearning. For AMUN, this means that $\mathcal{D_{\text{A}}}$ is computed on an updated model after each set of unlearning requests (shown as AMUN-A). In addition to comparing AMUN-A to the other unlearning methods, we also compare it to a version (shown as AMUN) that computes all the adversarial examples on the original model so it can handle the unlearning requests faster upon receiving them i.e., $\mathcal{D_{\text{A}}}$ is not computed on an updated model after each request; the set of requests are batched and $\mathcal{D_{\text{A}}}$ is computed on the entire batch. For this experiment, we use a ResNet-18 model trained on training set of CIFAR-10 ($50$K samples). Our goal is to unlearn $10\%$ of the training samples ($5$K), but this time in $5$ consecutive sets of size $2\%$ ($1$K) each. We then evaluate the effectiveness of unlearning at each step using RMIA.
@@ 8 Acknowledgment
This work used Delta computing resources at National Center for Supercomputing Applications through allocation CIS240316 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program Boerner et al. (2023), which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
@@ Appendix A Proofs
where the inequality is derived by using the smoothness property (^$\|\nabla\hat{\mathcal{R}}(\theta_{o})\|_{2}^{2}\leq 2\beta(\hat{\mathcal{R}}(\theta_{o})-\hat{\mathcal{R}}(\theta^{\prime}))$) and the convexity assumption which leads to the inequality: ^$\hat{\mathcal{R}}(\theta_{o}))\geq\hat{\mathcal{R}}(\theta_{u})+\langle\nabla\hat{\mathcal{R}}(\theta_{o}),\theta_{o}-\theta_{u}\rangle$.
Next, we derive an upper-bound for ^$\hat{\mathcal{R}}(\theta_{u})-\hat{\mathcal{R}}(\theta^{\prime})$ to replace in the above inequality. By the definition of unnormalized empirical loss on $\mathcal{D}^{\prime}$:
^$\hat{\mathcal{R}}(\theta_{u})-\hat{\mathcal{R}}(\theta^{\prime})$
@@
where the last equality was derived by the assumption that models are trained until they achieve near-00 loss on their corresponding dataset. Therefore, $\sum_{i=1}^{n-1}\ell(f_{\theta_{u}}(x_{i}),y_{i})=\sum_{i=1}^{n-1}\ell(f_{\theta^{\prime}}(x_{i}),y_{i})=0$ since the retrained model has been trained on the remaining samples and the unlearned model has been derived by a single step of gradient descent on the original model, that had been trained on $\mathcal{D}$.
To further simplify the derived terms above and reaching at our desired inequality, we focus on the term $-\ell(f_{\theta^{\prime}}(x),y)$. By adding and decreasing the term $\ell(f_{\theta_{o}}(x^{\prime}),y)$ we get:
@@
To improve upon fine-tuning on samples in $\mathcal{D_{\text{F}}}$ with randomly chosen wrong labels, Chen et al. [2023b] use the labels derived from one step of the FGSM attack to choose the new labels for the samples in $\mathcal{D_{\text{F}}}$. This method which was presented as BS in our experiments (§ 6.1), does not use the adversarial examples and only uses their labels as the new labels for samples of $\mathcal{D_{\text{F}}}$. This corresponds to the dataset $\mathcal{D}_{AdvL}$ in § 6.2.1. As our results in Figures 1 and 5 show, fine-tuning the trained model on this dataset leads to catastrophic forgetting even when $\mathcal{D_{\text{R}}}$ is available. This is simply due to the fact that the samples in $\mathcal{D}_{AdvL}$ contradict the distribution that the trained models have already learned.
The work by Setlur et al. [2022] is not an unlearning method, despite what the name suggest. They propose a regularization method that tries to maximize the loss on the adversarial examples of the training samples that are relatively at a higher distance to lower the confidence of the model on those examples. The work by Zhang et al. [2024b] proposed a defense method similar to adversarial training for making to unlearned LLMs more robust to jailbreak attacks on the topics that they have unlearned. Łucki et al. [2024] also study the careful application of jailbreak attacks against unlearned models. The work by Jung et al. [2024] investigate computing adversarial noise to mask the model parameters. Many of the works with similar titles, use “adversarial" to refer to minimax optimization [Zeng et al., 2021] or considering a Stackelberg game setting between the source model and the adversary that is trying to extract information [Di et al., 2024].
@@ Appendix C Implementation Details
For all the experiments we train three models on $\mathcal{D}$. For each size of $\mathcal{D}$ ($10\%$ or $50\%$), we use three random subsets and for each subset, we try three different runs of each of the unlearning methods. This leads to a total of $27$ runs of each unlearning method using different initial models and subsets of $\mathcal{D}$ to unlearn. Hyper-parameter tuning of each of the methods is done on a separate random subset of the same size from $\mathcal{D}$, and then the average performance is computed for the other random subsets used as $\mathcal{D_{\text{F}}}$. For tuning the hyper-parameters of the models, we followed the same range suggested by their authors and what has been used in the prior works for comparisons. Similar to prior works Liu et al. [2024], Fan et al. [2024], we performed 10 epochs for each of the unlearning methods, and searched for best learning rate and number of steps for a learning rate scheduler. More specifically, for each unlearning method, we performed a grid search on learning rates within the range of $[10^{-6},10^{-1}]$ with an optional scheduler that scales the learning rate by $0.1$ for every $1$ or $5$ steps. For SalUn, whether it is used on its own or in combination with AMUN, we searched for the masking ratios in the range $[0.1,0.9]$.
@@ Appendix D AMUN + SalUn
The main idea behind SalUn is to limit the fine-tuning of the model, during unlearning, to only a subset of the parameters of the model, while keeping the rest of them fixed. Fan et al. [2024] show that this technique helps to preserve the accuracy of the model when fine-tuning the model on $\mathcal{D_{\text{F}}}$ with randomly-chosen wrong labels. More specifically, they compute a mask using the following equation:
@@
[…]
@@ E.1 Empirical Behavior of Retrained Models
where $f(x)_{y}$ is the predicted probability for the correct class. This scaling has been used by Carlini et al. [2022] to transform the the prediction probabilities such that they can be better approximated with a normal distribution, which are indeed used by some of the SOTA MIA methods for predicting training samples from the test samples Carlini et al. [2022]. Figure 3 (left) shows these fitted normal distribution as well, which perfectly match for $\mathcal{D_{\text{F}}}$ and $\mathcal{D_{\text{T}}}$.
@@ E.1.1 Confidence Values in Unlearned Models
We also evaluated AMUN for unlearning in models that are adversarially trained. we performed our analysis on ResNet-18 models trained using TRADES loss Zhang et al. [2019] on CIFAR-10. We performed the experiments for unlearning 10% of the dataset in both cases where $\mathcal{D_{\text{R}}}$ is accessible and not. As the results in Table LABEL: show, in both settings AMUN is effective in unlearning the forget samples and achieving a low gap with the retrained models. This gap is obviously smaller when there is access to $\mathcal{D_{\text{R}}}$.
@@ E.4 Transferred Adversarial Examples
One of the intriguing properties of adversarial attacks is their transferability to other models Papernot et al. [2016], Liu et al. [2016]; Adversarial examples generated on a trained model (source model) mostly transfer successfully to other models (target models). This success rate of the transferred adversarial examples increases if the source model and target model have the same architecture Papernot et al. [2016]. There are other studies that can be used to increase the success rate of this type of attack Zhao et al. [2021], Zhang et al. [2022], Chen et al. [2023a], Ebrahimpour-Boroojeny et al. [2024]. In this section, we are interested to see if using the the adversarial examples generated using Algorithm 1 for a given model trained on some dataset $\mathcal{D}$ can be used as the $\mathcal{D_{\text{A}}}$ dataset for unlearning a portion of $\mathcal{D}$ from a separately trained model. The advantage of using adversarial examples generated for another model is saving the computation cost for other trained models. For this purpose, we train three ResNet-18 models separately on CIFAR-10, we generate the adversarial examples for each of these models using Algorithm 1. We use AMUN for unlearning $10\%$ and $50\%$ of CIFAR-10 from either of these models, but instead of their adversarial samples, we use the ones derived from the other models. The results in Table 6 shows that using transferred adversarial examples leads to lower performance, specially for the case where there is no access to $\mathcal{D_{\text{R}}}$. All the values for test accuracy are also lower compared to using adversarial examples from the model itself because these adversarial examples from the other models do not all belong to the natural distribution of the model and they do not even always transfer to the other models. Still the results are comparable to the prior SOTA methods in unlearning, and even in the case of no access to $\mathcal{D_{\text{R}}}$ outperforms all prior methods.
@@ E.5 Weak Attacks
In this section we evaluate the effectiveness of using weaker attacks in Algorithm 1. For this purpose, we perform the unlearning on a ResNet-18 model trained on CIFAR-10 in all unlearning settings mentioned in § 5, and compare the results with the default choice of PGD-50 in AMUN. The weaker attack that we use is a variation of FFGSM Wong et al. [2020], which itself is a variant of FGSM Goodfellow et al. [2014]. FGSM takes steps toward the gradient sign at a given sample to find adversarial samples. FFGSM takes a small step toward a random direction first, and then proceeds with FGSM. To adapt these method to the format of Algorithm 1 we start with FGSM attack; we find the gradient sign and start to move toward that direction in steps of size $\epsilon$ until we find an adversarial example. If the adversarial example is not found after a few iteration of the While loop, we restart the value of $\epsilon$ and add a small random perturbation before the next round of FGSM attack and the While loop. We continue this procedure to find an adversarial sample. After deriving a new set of adversarial examples using this methods, we performed a separate round hyper-parameter tuning for unlearning with the new attack to have a fair comparison. It is notable to mention that this leads to a much faster attack because we only compute the gradient once for each round round of FGSM (at the beginning or after each addition of random perturbation and restarting FGSM). Table 7 shows the comparison of the results with the original version of AMUN that uses PGD-50. As the results show, using this weaker attack leads to worse results; however, they still outperform prior SOTA methods in unlearning, specially in the setting where there is no access to $\mathcal{D_{\text{R}}}$ and the size of $\mathcal{D_{\text{F}}}$ is $50\%$ of $\mathcal{D}$.

### v3

@@ 6.2.2 Adversarially Robust Models
We evaluate the effectiveness of AMUN when the trained model is adversarially robust. One of the most effective methods in designing robust models is adversarial training which targets smoothing the model’s prediction function around the training samples salman2019provably. This has been shown to provably enhance the adversarial robustness of the model cohen2019certified. One of the effective adversarial training methods is by using TRADES loss introduced by (zhang2019theoretically). We will use adversarially trained ResNet-18 models for unlearning $10\%$ of the samples in CIFAR-10. In addition, we will use another defense mechanism that is less costly and more practical for larger models. There is a separate line of work that try to achieve the same smoothness in model’s prediction boundary by controlling the Lipschitz constant of the models szegedy2013intriguing. The method proposed by boroojeny2024spectrum is much faster than adversarial training and their results show a significant improvement in the robust accuracy. We use their clipping method to evaluate the effectiveness of AMUN for unlearning $10\%$ and $50\%$ of the samples from robust ResNet-18 models trained on CIFAR-10.
Results: Table 5 in Appendix E.2 shows the results for the adversarially trained models. For the models with controlled Lipschitz continuity, the results are shows in Table 3 (no access to $\mathcal{D_{\text{R}}}$) and Table 4 in Appendix E.2 (with access to $\mathcal{D_{\text{R}}}$). As the results show, even when there is no access to $\mathcal{D_{\text{R}}}$, AMUN still results in effective unlearning for adversarially robust models; RMIA does not do better than random guessing for predicting $\mathcal{D_{\text{F}}}$ from $\mathcal{D_{\text{T}}}$. As Fig. 3 (right) shows, in the robust models, more than $97\%$ of the adversarial examples are further away from their corresponding training samples, compared to this distance for the original models. However, this does not interfere with the performance of AMUN because these robust models are smoother and tend to be more regularized. This regularization, which prevents them from overfitting to the training samples is in fact a contributing factor to the improved generalization bounds for these models bartlett2017spectrally. This in itself contributes to enhanced resilience against MIAs. As seen in Tables 3 and 4, even for the clipped models retrained on $\mathcal{D_{\text{R}}}$, the AUC score of RMIA for predicting $\mathcal{D_{\text{R}}}$ from $\mathcal{D_{\text{F}}}$ (FR AUC) is very low, which shows that these smoother models .
@@ 6.3 Continuous Unlearning
We evaluate the performance of the unlearning methods when they are used to perform multiple consecutive unlearning from a trained model. This is a desirable capability for unlearning methods because in real-world applications there might be multiple unlearning requests and it is preferred to minimize the number of times that a model needs to be retrained from scratch. The setting we envision is as follows: models are updated at each request for unlearning. For AMUN, this means that $\mathcal{D_{\text{A}}}$ is computed on an updated model after each set of unlearning requests (shown as AMUN-A). In addition to comparing AMUN-A to the other unlearning methods, we also compare it to a version (shown as AMUN) that computes all the adversarial examples on the original model so it can handle the unlearning requests faster upon receiving them i.e., $\mathcal{D_{\text{A}}}$ is not computed on an updated model after each request; the set of requests are batched and $\mathcal{D_{\text{A}}}$ is computed on the entire batch. For this experiment, we use a ResNet-18 model trained on training set of CIFAR-10 (5050K samples). Our goal is to unlearn $10\%$ of the training samples (55K), but this time in $5$ consecutive sets of size $2\%$ (11K) each. We then evaluate the effectiveness of unlearning at each step using RMIA.
@@ 8 Acknowledgment
This work used Delta computing resources at National Center for Supercomputing Applications through allocation CIS240316 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program boerner2023access, which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
@@ Appendix A Proofs
where the inequality is derived by using the smoothness property ($\|\nabla\hat{\mathcal{R}}(\theta_{o})\|_{2}^{2}\leq 2\beta(\hat{\mathcal{R}}(\theta_{o})-\hat{\mathcal{R}}(\theta^{\prime}))$) and the convexity assumption which leads to the inequality: $\hat{\mathcal{R}}(\theta_{o}))\geq\hat{\mathcal{R}}(\theta_{u})+\langle\nabla\hat{\mathcal{R}}(\theta_{o}),\theta_{o}-\theta_{u}\rangle$.
Next, we derive an upper-bound for $\hat{\mathcal{R}}(\theta_{u})-\hat{\mathcal{R}}(\theta^{\prime})$ to replace in the above inequality. By the definition of unnormalized empirical loss on $\mathcal{D}^{\prime}$:
$\hat{\mathcal{R}}(\theta_{u})-\hat{\mathcal{R}}(\theta^{\prime})$
@@
where the last equality was derived by the assumption that models are trained until they achieve near-0 loss on their corresponding dataset. Therefore, $\sum_{i=1}^{n-1}\ell(f_{\theta_{u}}(x_{i}),y_{i})=\sum_{i=1}^{n-1}\ell(f_{\theta^{\prime}}(x_{i}),y_{i})=0$ since the retrained model has been trained on the remaining samples and the unlearned model has been derived by a single step of gradient descent on the original model, that had been trained on $\mathcal{D}$.
To further simplify the derived terms above and reaching at our desired inequality, we focus on the term $\ell(f_{\theta^{\prime}}(x),y)$. By adding and decreasing the term $\ell(f_{\theta_{o}}(x^{\prime}),y)$ we get:
@@
To improve upon fine-tuning on samples in $\mathcal{D_{\text{F}}}$ with randomly chosen wrong labels, chen2023boundary use the labels derived from one step of the FGSM attack to choose the new labels for the samples in $\mathcal{D_{\text{F}}}$. This method which was presented as BS in our experiments (§ 6.1), does not use the adversarial examples and only uses their labels as the new labels for samples of $\mathcal{D_{\text{F}}}$. This corresponds to the dataset $\mathcal{D}_{AdvL}$ in § 6.2.1. As our results in Figures 1 and 5 show, fine-tuning the trained model on this dataset leads to catastrophic forgetting even when $\mathcal{D_{\text{R}}}$ is available. This is simply due to the fact that the samples in $\mathcal{D}_{AdvL}$ contradict the distribution that the trained models have already learned.
The work by setlur2022adversarial is not an unlearning method, despite what the name suggest. They propose a regularization method that tries to maximize the loss on the adversarial examples of the training samples that are relatively at a higher distance to lower the confidence of the model on those examples. The work by zhang2024defensive proposed a defense method similar to adversarial training for making to unlearned LLMs more robust to jailbreak attacks on the topics that they have unlearned. lucki2024adversarial also study the careful application of jailbreak attacks against unlearned models. The work by jung2024attack investigate computing adversarial noise to mask the model parameters. Many of the works with similar titles, use “adversarial" to refer to minimax optimization [zeng2021adversarial] or considering a Stackelberg game setting between the source model and the adversary that is trying to extract information [di2024adversarial].
@@ Appendix C Implementation Details
For all the experiments we train three models on $\mathcal{D}$. For each size of $\mathcal{D}$ ($10\%$ or $50\%$), we use three random subsets and for each subset, we try three different runs of each of the unlearning methods. This leads to a total of $27$ runs of each unlearning method using different initial models and subsets of $\mathcal{D}$ to unlearn. Hyper-parameter tuning of each of the methods is done on a separate random subset of the same size from $\mathcal{D}$, and then the average performance is computed for the other random subsets used as $\mathcal{D_{\text{F}}}$. For tuning the hyper-parameters of the models, we followed the same range suggested by their authors and what has been used in the prior works for comparisons. Similar to prior works liu2024model, fan2023salun, we performed 10 epochs for each of the unlearning methods, and searched for best learning rate and number of steps for a learning rate scheduler. More specifically, for each unlearning method, we performed a grid search on learning rates within the range of $10^{-6},10^{-1}]$ with an optional scheduler that scales the learning rate by $0.1$ for every $1$ or $5$ steps. For SalUn, whether it is used on its own or in combination with AMUN, we searched for the masking ratios in the range $[0.1,0.9]$.
@@ Appendix D AMUN + SalUn
The main idea behind SalUn is to limit the fine-tuning of the model, during unlearning, to only a subset of the parameters of the model, while keeping the rest of them fixed. fan2023salun show that this technique helps to preserve the accuracy of the model when fine-tuning the model on $\mathcal{D_{\text{F}}}$ with randomly-chosen wrong labels. More specifically, they compute a mask using the following equation:
@@
Figure 3: (left) These two plots show the histogram of confidence values of the retrained model on its predictions for the remaining set (Remain), test set (Test), and forget set (Forget) during the training, when the size of the forget set is $\%10$ (1st plot) and $\%50$ (2nd plot) of the training set. It also shows the Gaussian distributions fitted to each histogram. As the plots show the models perform similarly on the forget set and test set because to the retrained model they are unseen data from the same distribution. (right) This plot compares the $\delta_{x}$ value in definition 2.1 for adversarial examples generated on the original ResNet-18 models (x-axis) and clipped ResNet-18 models (y-axis). The dashed line shows $x=y$ line and more than $97\%$ of the values fall bellow this line.
@@ E.1 Empirical Behavior of Retrained Models
where $f(x)_{y}$ is the predicted probability for the correct class. This scaling has been used by carlini2022membership to transform the the prediction probabilities such that they can be better approximated with a normal distribution, which are indeed used by some of the SOTA MIA methods for predicting training samples from the test samples carlini2022membership. Figure 3 (left) shows these fitted normal distribution as well, which perfectly match for $\mathcal{D_{\text{F}}}$ and $\mathcal{D_{\text{T}}}$.
@@ E.1.1 Confidence Values in Unlearned Models
We also evaluated AMUN for unlearning in models that are adversarially trained. we performed our analysis on ResNet-18 models trained using TRADES loss zhang2019theoretically on CIFAR-10. We performed the experiments for unlearning 10% of the dataset in both cases where $\mathcal{D_{\text{R}}}$ is accessible and not. As the results in Table LABEL: show, in both settings AMUN is effective in unlearning the forget samples and achieving a low gap with the retrained models. This gap is obviously smaller when there is access to $\mathcal{D_{\text{R}}}$.
@@ E.4 Transferred Adversarial Examples
One of the intriguing properties of adversarial attacks is their transferability to other models papernot2016transferability, liu2016delving; Adversarial examples generated on a trained model (source model) mostly transfer successfully to other models (target models). This success rate of the transferred adversarial examples increases if the source model and target model have the same architecture papernot2016transferability. There are other studies that can be used to increase the success rate of this type of attack zhao2021success, zhang2022improving, chen2023rethinking, ebrahimpour2024lotos. In this section, we are interested to see if using the the adversarial examples generated using Algorithm 1 for a given model trained on some dataset $\mathcal{D}$ can be used as the $\mathcal{D_{\text{A}}}$ dataset for unlearning a portion of $\mathcal{D}$ from a separately trained model. The advantage of using adversarial examples generated for another model is saving the computation cost for other trained models. For this purpose, we train three ResNet-18 models separately on CIFAR-10, we generate the adversarial examples for each of these models using Algorithm 1. We use AMUN for unlearning $10\%$ and $50\%$ of CIFAR-10 from either of these models, but instead of their adversarial samples, we use the ones derived from the other models. The results in Table 6 shows that using transferred adversarial examples leads to lower performance, specially for the case where there is no access to $\mathcal{D_{\text{R}}}$. All the values for test accuracy are also lower compared to using adversarial examples from the model itself because these adversarial examples from the other models do not all belong to the natural distribution of the model and they do not even always transfer to the other models. Still the results are comparable to the prior SOTA methods in unlearning, and even in the case of no access to $\mathcal{D_{\text{R}}}$ outperforms all prior methods.
@@ E.5 Weak Attacks
In this section we evaluate the effectiveness of using weaker attacks in Algorithm 1. For this purpose, we perform the unlearning on a ResNet-18 model trained on CIFAR-10 in all unlearning settings mentioned in § 5, and compare the results with the default choice of PGD-50 in AMUN. The weaker attack that we use is a variation of FFGSM wong2020fast, which itself is a variant of FGSM goodfellow2014explaining. FGSM takes steps toward the gradient sign at a given sample to find adversarial samples. FFGSM takes a small step toward a random direction first, and then proceeds with FGSM. To adapt these method to the format of Algorithm 1 we start with FGSM attack; we find the gradient sign and start to move toward that direction in steps of size $\epsilon$ until we find an adversarial example. If the adversarial example is not found after a few iteration of the While loop, we restart the value of $\epsilon$ and add a small random perturbation before the next round of FGSM attack and the While loop. We continue this procedure to find an adversarial sample. After deriving a new set of adversarial examples using this methods, we performed a separate round hyper-parameter tuning for unlearning with the new attack to have a fair comparison. It is notable to mention that this leads to a much faster attack because we only compute the gradient once for each round round of FGSM (at the beginning or after each addition of random perturbation and restarting FGSM). Table 7 shows the comparison of the results with the original version of AMUN that uses PGD-50. As the results show, using this weaker attack leads to worse results; however, they still outperform prior SOTA methods in unlearning, specially in the setting where there is no access to $\mathcal{D_{\text{R}}}$ and the size of $\mathcal{D_{\text{F}}}$ is $50\%$ of $\mathcal{D}$.